# Unifying the Three Algebraic Approaches to the CSP via Minimal Taylor Algebras

### v5

@@ 7.5 Absorption and pp-definitions
Pairs of subpowers of $\mathbf{A}$ with $\mathbf{B}\trianglelefteq\mathbf{C}$ are also closed under pp-definitions in the following sense (this folklore fact was recorded as Lemma 2.9 in  [3]).
@@
where $R_{1},\dots,R_{k}$ are subpowers of $\mathbf{A}$ regarded as predicates and $\sigma_{1}$,…,$\sigma_{k}$ stand for sequences of (free or bound) variables. Let $S_{1},\dots,S_{k}$ be subpowers of $\mathbf{A}$ such that $S_{i}\trianglelefteq\mathbf{R}_{i}$ for all $i$. Then the subpower
@@
absorbs $\mathbf{R}$. Moreover, if all the absorptions $S_{i}\trianglelefteq\mathbf{R}_{i}$ are witnessed by $t$, then so is $S\trianglelefteq\mathbf{R}$.
In particular, if $R\leq_{sd}\mathbf{A}^{2}$ and $B$ absorbs $\mathbf{A}$, then $B+R$ and $B-R$ absorb $\mathbf{A}$ as well.
It is also useful to observe that absorption is transitive: if $\mathbf{C}\trianglelefteq\mathbf{B}$ by $s$ and $\mathbf{B}\trianglelefteq\mathbf{A}$ by $t$, then $\mathbf{C}\trianglelefteq\mathbf{A}$ by the star composition of $s$ and $t$.
@@ 7.6 Relational descriptions
The first one is a characterization of absorption by means of so called $B$-essential relations (see e.g. Proposition 2.14 in [3]). A relation $R\subseteq A^{n}$ is $B$-essential if $R$ does not intersect $B^{n}$ but every projection of $R$ onto all but one of the coordinates intersects the corresponding power of $B$. The characterization is that $B\trianglelefteq_{n}\mathbf{A}$ if, and only if, there are no $B$-essential subuniverses of $\mathbf{A}^{n}$. We state this fact as follows.
@@
Let $\mathbf{A}$ be an algebra and $B\leq\mathbf{A}$. Then $B\trianglelefteq_{n}\mathbf{A}$ if and only if for every ${\mathbf{a}}^{1},\dotsc,{\mathbf{a}}^{n}\in A^{n}$ such that ${\mathbf{a}}^{i}_{j}\in B$ for $i\neq j$ we have $\operatorname{Sg}_{\mathbf{A}^{n}}({\mathbf{a}}^{1},\dotsc,{\mathbf{a}}^{n})\cap B^{n}\neq\emptyset$.
@@
Let $\mathbf{A}$ be an algebra and $B\subseteq A$. Then $B$ is a strongly projective subuniverse of $\mathbf{A}$ if and only if the relation $R(x,y,z)=B(x)\vee(y=z)$ is a subuniverse of $\mathbf{A}^{3}$.
Proof.
For the backward implication let $f$ be an $n$-ary term operation of $\mathbf{A}$ and say, without loss of generality, that the first coordinate is essential as witnessed by tuples $(c,c_{2},\dotsc,c_{n})$ and $(c^{\prime},c_{2},\dotsc,c_{n})$. Take $(b,a_{2},\dotsc,a_{n})\in B\times A^{n-1}$ and note that $R(b,c,c^{\prime})$, $R(a_{2},c_{2},c_{2}),\dotsc,R(a_{n},c_{n}$, $c_{n})$. Therefore
@@
For the forward implication we proceed by way of contradiction and suppose that an application of an operation $f$ to triples from $R$ produces a triple outside. The resulting triple does not have an element of $B$ at the first position, therefore, by the assumption, all the input triples that have an element of $B$ on the first position appear on inessential coordinates of $f$. The remaining triples have the same element on the second and third positions, therefore so does the resulting triple, a contradiction. ∎
@@ 8.1 Absorption theorem
The first step toward Corollary 4.5 can be divided into two sub-steps. The first sub-step is Lemma 2.7 in [8].
@@
The other sub-step has a short proof via Taylor operations, see the proof of Lemma 3.4 in [8]. An elementary proof via clone homomorphisms is also available, but unpublished. It will appear in another paper.
@@
For any set $D=\{c_{1},\dots,c_{k}\}\subseteq C$ consider the binary pp-definable relation $S_{D}(x,y)$ expressing “$x$ , $y$, and all the $c_{i}$ have a common neighbor”, that is $S_{D}(x,y)\equiv(\exists a)R(a,x)\wedge R(a,y)\wedge R(a,c_{1})\wedge\dots\wedge R(a,c_{k})$. Since $-R+R=C^{2}$, the set $S_{\emptyset}$ is equal to $C^{2}$, and since $R$ is not left central, the set $S_{C}$ is empty. Take a maximal $D$ so that $S_{D}=C^{2}$ and take any $E\subseteq C$ with $D\subseteq E$ and $|E\setminus D|=1$. Now observe that $S_{E}$ is a proper symmetric subdirect relation on $C$ whose center is nonempty since it contains $E$. ∎
The third step shows how left central relations together with transitive operations on the right side produce absorption. The argument is in the final part of the proof of Theorem 2.11 in [4].
@@
Let $\mathbf{A}$ and $\mathbf{C}$ be algebras and suppose $R\leq_{sd}\mathbf{A}\times\mathbf{C}$ is left-central. If $\mathbf{C}$ has a transitive term operation $t^{\mathbf{C}}$, then the left center of $R$ absorbs $\mathbf{A}$ by $t^{\mathbf{A}}$.
Proof.
Let $t^{\mathbf{C}}$ be a transitive term operation of $\mathbf{C}$, say of arity $n$, and let $B$ be the left center of $R$. We need to show that, for any ${\mathbf{a}}\in A^{n}$ such that |$|\{i:a_{i}\in B\}|\geq n-1$, the result $t^{\mathbf{A}}({\mathbf{a}})=b$ is in $B$, i.e., we need to show that for every $c\in C$ we have $(b,c)\in R$. Let $i$ be the only coordinate for which $a_{i}\not\in B$ (or take $i$ arbitrary if there is none such) and take $d$ such that $(a_{i},d)\in R$. By transitivity of $t^{\mathbf{C}}$, there exists ${\mathbf{d}}\in C^{n}$ with $d_{i}=d$ and $t^{\mathbf{C}}({\mathbf{d}})=c$. All the component pairs $(a_{j},d_{j})$ are in $R$ since $B$ is the left center of $R$, so $(t^{\mathbf{A}}({\mathbf{a}}),t^{\mathbf{C}}({\mathbf{d}}))=(b,c)$ is in $R$, as required. ∎
Note that the three steps combined already give us the absorption theorem (Theorem 3.4). We prove a slightly stronger version where only $\mathbf{C}$ is assumed to be Taylor.
@@
Suppose $R\leq_{sd}\mathbf{A}\times\mathbf{C}$ is proper and linked, and $\mathbf{C}$ is Taylor. Then $\mathbf{A}$ or $\mathbf{C}$ has a nontrivial absorbing subuniverse.
Proof.
Proposition 4.1 implies that there is a left central subuniverse $S$ of $\mathbf{A}\times\mathbf{C}$ or $\mathbf{C}^{2}$. The left center $B$ of $S$ (which is a subuniverse of $\mathbf{A}$ or $\mathbf{C}$ depending on the case) is a Taylor center unless $\mathbf{C}$ contains a nontrivial proper 2-absorbing subuniverse (in which case we are done). Then $\mathbf{C}$ has no nontrivial projective subuniverses by Proposition 8.2 and therefore it has a transitive term operation by Proposition 8.1. Proposition 8.3 finishes the proof by showing that $B$ is an absorbing subuniverse (of $\mathbf{A}$ or $\mathbf{C}$). ∎
@@
Suppose $R\leq_{sd}\mathbf{A}\times\mathbf{C}$ is left-central and $\mathbf{C}$ is Taylor. Then $\mathbf{C}$ has a nontrivial projective and 2-absorbing subuniverse, or the left center of $R$ absorbs $\mathbf{A}$.
Proof.
Either $\mathbf{C}$ has a nontrivial projective subuniverse (which is 2-absorbing by Proposition 8.2) or $\mathbf{C}$ has a transitive term operation by Proposition 8.1 and then the left center absorbs $\mathbf{A}$ by Proposition 8.3. ∎
Now we improve absorption to 3-absorption. We further divide this task into two sub-steps. The first one isolates a property of centers that was, inspired by this paper, exploited in [61].
@@
Let $B$ be a center of $\mathbf{A}$ and let $a\in A\setminus B$. Then $(a,a)$ is not in the subuniverse of $\mathbf{A}^{2}$ generated by ($(\{a\}\times B)\cup(B\times B)\cup(B\times\{a\})$.
Proof.
Let $R\leq_{sd}\mathbf{A}\times\mathbf{C}$ be a witness of centrality. Suppose, for a contradiction that $(a,a)$ is generated by a term operation $f^{\mathbf{A}}$, so
$f^{\mathbf{A}}(a,\dotsc,a,b_{1},\dotsc,b_{i})=a=f^{\mathbf{A}}(b^{\prime}_{1},\dotsc,b^{\prime}_{j},a,\dotsc,a)$
for some $b_{1},\dots,b_{i},b_{1}^{\prime},\dots,b_{j}^{\prime}\in B$ where $i+j$ is not less than the arity of $f$. Therefore $f^{\mathbf{C}}(a+R,\dots,a+R,b_{1}+R,\dots,b_{i}+R)\subseteq a+R$ and, denoting $D=a+R$, we have $f^{\mathbf{C}}(D,\dots,D,C,\dots,C)\subseteq D$ (with $i$ occurrences of $C$ on the left). Similarly, we obtain $f^{\mathbf{C}}(C,\dots,C,D,\dots,D)\subseteq D$ with $j$ occurrences of $C$. It follows that the binary operation on $C$ obtained from $f^{\mathbf{C}}$ by identifying the first $j$ variables to $x$ and the rest to $y$ witnesses the nontrivial 2-absorption $D\trianglelefteq_{2}\mathbf{C}$, a contradiction with the definition of a center. ∎
@@
Suppose $B\trianglelefteq\mathbf{A}$ and that for every $a\in A\setminus B$ we have $(a,a)\notin\operatorname{Sg}_{\mathbf{A}^{2}}(\{a\}\times B\cup B\times\{a\})$. Then $B$ $3$-absorbs $\mathbf{A}$.
Proof.
Let $n$ be the minimal number such that $B\trianglelefteq_{n+1}\mathbf{A}$ and assume, striving for a contradiction, that $n>2$. By Proposition 7.2 there exist ${\mathbf{a}}^{1},\dotsc,{\mathbf{a}}^{n}\in A^{n}$ such that $a^{i}_{j}\in B$ for $i\neq j$ and $\operatorname{Sg}_{\mathbf{A}^{n}}({\mathbf{a}}^{1},\dotsc,{\mathbf{a}}^{n})\cap B^{n}=\emptyset$. Put $R=\operatorname{Sg}_{\mathbf{A}^{n}}({\mathbf{a}}^{1},\dotsc,{\mathbf{a}}^{n})$ and assume that $R$ is an inclusion minimal relation among all choices of ${\mathbf{a}}^{1},\dotsc,{\mathbf{a}}^{n}\in A^{n}$.
@@
$\operatorname{Sg}_{\mathbf{A}^{2}}\left((\{a^{n}_{n}\}\times B)\cup(B\times\{a^{n}_{n}\})\right).$
@@
We pp-define $R^{\prime}\leq\mathbf{A}^{2n-2}$ by the formula
@@
For $i\in\{1,\dots,n\}$ by ${\mathbf{c}}^{i}$ we denote ${\mathbf{a}}^{i}$ with the last coordinate removed. By the definition of $R^{\prime}$ and $S$, we have $({\mathbf{c}}^{i},{\mathbf{c}}^{n}),({\mathbf{c}}^{n},{\mathbf{c}}^{i})\in R^{\prime}$ for every $i\in\{1,\dots,n-1\}$. Moreover, these $22n-$ 2 tuples satisfy the condition in the second part of Proposition 7.2. Therefore, if $R^{\prime}\cap B^{2n-2}=\emptyset$, then $B\not\trianglelefteq_{2n-2}\mathbf{A}$. But $B\trianglelefteq_{n+1}\mathbf{A}$ and $12n-2\geq n+$ 1, a contradiction.
Since $R^{\prime}\cap B^{2n-2}\neq\emptyset$, there exist ${\mathbf{d}},{\mathbf{d}}^{\prime}\in R\cap(B^{n-1}\times A)$ such that $(d_{n},d^{\prime}_{n})\in S$. Let $E$ be the projection of $R$ onto the last coordinate after fixing all the other coordinates to $B$, that is, $E=\operatorname{proj}_{n}(R\cap(B^{n-1}\times A))$. Since $R$ was chosen inclusion minimal, we get $\operatorname{proj}_{n}(R)\subseteq\operatorname{Sg}_{\mathbf{A}}(B\cup\{e\})$ for every $e\in E$, otherwise we could replace ${\mathbf{a}}^{n}$ by a tuple ${\mathbf{b}}\in R\cap(B^{n-1}\times\{e\})$ and get a $B$-essential relation properly contained in $R$.
Let $E^{\prime}=E+S$. Since $B\cup\{d^{\prime}_{n}\}\subseteq E^{\prime}$ (as $a^{n}_{n},d_{n}\in E$, $(d_{n},d^{\prime}_{n})\in S$, and $(a^{n}_{n},b)\in S$ for all $b\in B$), we have $E^{\prime}\supseteq\operatorname{proj}_{n}(R)$, in particular $a_{n}^{n}\in E^{\prime}$. Therefore $(e,a_{n}^{n})\in S$ for some $e\in E$. It follows that $E^{\prime\prime}=a_{n}^{n}-S$ contains $B\cup\{e\}$ and we get $E^{\prime\prime}\supseteq\operatorname{proj}_{n}(R)$, in particular $a_{n}^{n}\in E^{\prime\prime}$. So $(a^{n}_{n},a^{n}_{n})\in S$, which contradicts our assumption on $B$. ∎
The argument in Corollary 8.4 together with Lemma 8.6 and Proposition 8.7 now proves Proposition 4.3 (which is Corollary 7.10.2 in [60]) and also the improved absorption theorem.
@@
Proof.
The algebra $\mathbf{C}$ has no nontrivial projective subuniverses by Proposition 8.2 and therefore it has a transitive term operation by Proposition 8.1. By Proposition 8.3, $B$ is an absorbing subuniverse of $\mathbf{A}$. Lemma 8.6 and Proposition 8.7 now imply $B\trianglelefteq_{3}\mathbf{A}$. ∎
@@
Proof.
Proposition 4.1 implies that there is a left central subuniverse of $\mathbf{A}\times\mathbf{C}$ or $\mathbf{C}^{2}$. In both cases, Proposition 4.3 finishes the proof. ∎
@@
Proof.
Let $R\leq_{sd}\mathbf{A}\times\mathbf{C}$ be the witnessing relation from the definition of center. We define a directed graph on $A$, by putting $c\rightarrow d$ if there exists a cyclic term $t$ and a choice of elements $b_{2},\dotsc,b_{n}\in B$ such that $t^{\mathbf{A}}(c,b_{2},\dotsc,b_{n})=d$.
Obviously the graph has no sinks, but it also easy to see that it is transitively closed. Indeed, let $c\rightarrow c^{\prime}$ be witnessed by $t$ and $b_{2},\dotsc,b_{n}$, and $c^{\prime}\rightarrow c^{\prime\prime}$ be witnessed by $t^{\prime}$ and $b_{2}^{\prime},\dotsc,b_{n^{\prime}}^{\prime}$. Then the star composition of $t^{\prime}$ and $t$, together with the tuple
@@
Next we claim that there is no $a\in A\setminus B$ with a self-loop. For a contradiction, say $a$ is such and let $t^{\mathbf{A}}(a,b_{2},\dotsc,b_{n})=a$ with $a\in A\setminus B$ and $b_{2},\dotsc,b_{n}\in B$. The set $D=a+R$ is a subuniverse of $\mathbf{C}$, and moreover
$t^{\mathbf{C}}(a+R,b_{2}+R,\dotsc,b_{n}+R)=t^{\mathbf{C}}(D,C,\dotsc,C)\subseteq D$
and the same holds for all the cyclic shifts of $t$. Therefore $D$ is a $2$-absorbing subuniverse of $\mathbf{C}$ (by the operation $t^{\mathbf{C}}(x,y,\dots,y)$), a contradiction.
We conclude that all the loops and directed cycles in this graph must be inside $B$. Therefore there is a number $m$ so that starting with any $a\in A$, and following any directed walk, after $m$ steps we necessarily arrive in $B$. It suffices to take any cyclic term of $\mathbf{A}$ and star-compose it with itself $m$ times to obtain an operation which witnesses $\mathbf{B}\trianglelefteq\mathbf{A}$. Now it is enough to apply Lemma 8.6 and Proposition 8.7
@@
It follows that $lR\subseteq 2lR\subseteq 3lR\subseteq\ldots$, therefore this sequence of proper relations eventually stabilizes. We define $n=kl$ for a sufficiently large $k$ and obtain $2nR=nR$. ∎
@@
Thus $T_{j}$ is proper, subdirect, central and pp-definable from $R$. However, its left center is a proper superset of the left center of $R$ since it contains the element $a_{j}$ which is not in the left center of $R$. This is a contradiction with the choice of $R$. ∎

### v6

@@ 7.5 Absorption and pp-definitions
Pairs of subpowers of with are also closed under pp-definitions in the following sense (this folklore fact was recorded as Lemma 2.9 in  [Barto16:deciding_abs]).
@@
where $R_{1},\dots,R_{k}$ are subpowers of $\mathbf{A}$ regarded as predicates and $\sigma_{1}$,…,$\sigma_{k}$ stand for sequences of (free or bound) variables. Let $S_{1},\dots,S_{k}$ be subpowers of $\mathbf{A}$ such that $S_{i}\trianglelefteq$ for all $i$. Then the subpower
@@
absorbs . Moreover, if all the absorptions $S_{i}\trianglelefteq$ are witnessed by $t$, then so is $S\trianglelefteq$.
In particular, if $R\leq_{sd}$ and $B$ absorbs , then $B+R$ and $B-R$ absorb as well.
It is also useful to observe that absorption is transitive: if by $s$ and by $t$, then by the star composition of $s$ and $t$.
@@ 7.6 Relational descriptions
The first one is a characterization of absorption by means of so called $B$-essential relations (see e.g. Proposition 2.14 in [Barto16:deciding_abs]). A relation $R\subseteq A^{n}$ is $B$-essential if $R$ does not intersect $B^{n}$ but every projection of $R$ onto all but one of the coordinates intersects the corresponding power of $B$. The characterization is that $B\trianglelefteq_{n}$ if, and only if, there are no $B$-essential subuniverses of . We state this fact as follows.
@@
Let be an algebra and $B\leq$. Then $B\trianglelefteq_{n}$ if and only if for every ${\mathbf{a}}^{1},\dotsc,{\mathbf{a}}^{n}\in A^{n}$ such that ${\mathbf{a}}^{i}_{j}\in B$ for $i\neq j$ we have $\operatorname{Sg}({\mathbf{a}}^{1},\dotsc,{\mathbf{a}}^{n})\cap B^{n}\neq\emptyset$.
@@
Let be an algebra and $B\subseteq A$. Then $B$ is a strongly projective subuniverse of if and only if the relation $R(x,y,z)=B(x)\vee(y=z)$ is a subuniverse of $\mathbf{A}^{3}$.
Proof 7.4.
For the backward implication let $f$ be an $n$-ary term operation of and say, without loss of generality, that the first coordinate is essential as witnessed by tuples $(c,c_{2},\dotsc,c_{n})$ and $(c^{\prime},c_{2},\dotsc,c_{n})$. Take $(b,a_{2},\dotsc,a_{n})\in B\times A^{n-1}$ and note that $R(b,c,c^{\prime})$, $R(a_{2},c_{2},c_{2}),\dotsc,R(a_{n},c_{n}$, $c_{n})$. Therefore
@@
For the forward implication we proceed by way of contradiction and suppose that an application of an operation $f$ to triples from $R$ produces a triple outside. The resulting triple does not have an element of $B$ at the first position, therefore, by the assumption, all the input triples that have an element of $B$ on the first position appear on inessential coordinates of $f$. The remaining triples have the same element on the second and third positions, therefore so does the resulting triple, a contradiction.
@@ 8.1 Absorption theorem
The first step toward Corollary 4.5 can be divided into two sub-steps. The first sub-step is Lemma 2.7 in [Barto15:malcev].
@@
The other sub-step has a short proof via Taylor operations, see the proof of Lemma 3.4 in [Barto15:malcev]. An elementary proof via clone homomorphisms is also available, but unpublished. It will appear in another paper.
@@
For any set $D=\{c_{1},\dots,c_{k}\}\subseteq C$ consider the binary pp-definable relation $S_{D}(x,y)$ expressing “$x$ , $y$, and all the $c_{i}$ have a common neighbor”, that is $S_{D}(x,y)\equiv(\exists a)R(a,x)\wedge R(a,y)\wedge R(a,c_{1})\wedge\dots\wedge R(a,c_{k})$. Since $-R+R=C^{2}$, the set $S_{\emptyset}$ is equal to $C^{2}$, and since $R$ is not left central, the set $S_{C}$ is empty. Take a maximal $D$ so that $S_{D}=C^{2}$ and take any $E\subseteq C$ with $D\subseteq E$ and $|E\setminus D|=1$. Now observe that $S_{E}$ is a proper symmetric subdirect relation on $C$ whose center is nonempty since it contains $E$.
The third step shows how left central relations together with transitive operations on the right side produce absorption. The argument is in the final part of the proof of Theorem 2.11 in [Barto12:cyclic].
@@
Let and be algebras and suppose $R\leq_{sd}$ is left-central. If has a transitive term operation $t$, then the left center of $R$ absorbs by $t$.
Proof 8.5.
Let $t$ be a transitive term operation of , say of arity $n$, and let $B$ be the left center of $R$. We need to show that, for any ${\mathbf{a}}\in A^{n}$ such that |$|\{i:a_{i}\in B\}|\geq n-1$, the result $t({\mathbf{a}})=b$ is in $B$, i.e., we need to show that for every $c\in C$ we have $(b,c)\in R$. Let $i$ be the only coordinate for which $a_{i}\not\in B$ (or take $i$ arbitrary if there is none such) and take $d$ such that $(a_{i},d)\in R$. By transitivity of $t$, there exists ${\mathbf{d}}\in C^{n}$ with $d_{i}=d$ and $t({\mathbf{d}})=c$. All the component pairs $(a_{j},d_{j})$ are in $R$ since $B$ is the left center of $R$, so $(t({\mathbf{a}}),t({\mathbf{d}}))=(b,c)$ is in $R$, as required.
Note that the three steps combined already give us the absorption theorem (Theorem 3.4). We prove a slightly stronger version where only is assumed to be Taylor.
@@
Suppose $R\leq_{sd}$ is proper and linked, and is Taylor. Then or has a nontrivial absorbing subuniverse.
Proof 8.7.
Proposition 4.1 implies that there is a left central subuniverse $S$ of or . The left center $B$ of $S$ (which is a subuniverse of or depending on the case) is a Taylor center unless contains a nontrivial proper 2-absorbing subuniverse (in which case we are done). Then has no nontrivial projective subuniverses by Proposition 8.2, and therefore it has a transitive term operation by Proposition 8.1. Proposition 8.4 finishes the proof by showing that $B$ is an absorbing subuniverse (of or ).
@@
Suppose $R\leq_{sd}$ is left-central and is Taylor. Then has a nontrivial projective and 2-absorbing subuniverse, or the left center of $R$ absorbs .
Proof 8.9.
Either has a nontrivial projective subuniverse (which is 2-absorbing by Proposition 8.2) or has a transitive term operation by Proposition 8.1 and then the left center absorbs  by Proposition 8.4.
Now we improve absorption to 3-absorption. We further divide this task into two sub-steps. The first one isolates a property of centers that was, inspired by this paper, exploited in [zhuk21:strong-subalgs].
@@
Let $B$ be a center of and let $a\in A\setminus B$. Then $(a,a)$ is not in the subuniverse of generated by ($(\{a\}\times B)\cup(B\times B)\cup(B\times\{a\})$.
Proof 8.11.
Let $R\leq_{sd}$ be a witness of centrality. Suppose, for a contradiction that $(a,a)$ is generated by a term operation $f$, so
$f(a,\dotsc,a,b_{1},\dotsc,b_{i})=a=f(b^{\prime}_{1},\dotsc,b^{\prime}_{j},a,\dotsc,a)$
for some $b_{1},\dots,b_{i},b_{1}^{\prime},\dots,b_{j}^{\prime}\in B$ where $i+j$ is not less than the arity of $f$. Therefore $f(a+R,\dots,a+R,b_{1}+R,\dots,b_{i}+R)\subseteq a+R$ and, denoting $D=a+R$, we have $f(D,\dots,D,C,\dots,C)\subseteq D$ (with $i$ occurrences of $C$ on the left). Similarly, we obtain $f(C,\dots,C,D,\dots,D)\subseteq D$ with $j$ occurrences of $C$. It follows that the binary operation on $C$ obtained from $f$ by identifying the first $j$ variables to $x$ and the rest to $y$ witnesses the nontrivial 2-absorption $D\trianglelefteq_{2}$, a contradiction with the definition of a center.
@@
Suppose $B\trianglelefteq$ and that for every $a\in A\setminus B$ we have $(a,a)\notin\operatorname{Sg}(\{a\}\times B\cup B\times\{a\})$. Then $B$ $3$-absorbs .
Proof 8.13.
Let $n$ be the minimal number such that $B\trianglelefteq_{n+1}$ and assume, striving for a contradiction, that $n>2$. By Proposition 7.2 there exist ${\mathbf{a}}^{1},\dotsc,{\mathbf{a}}^{n}\in A^{n}$ such that $a^{i}_{j}\in B$ for $i\neq j$ and $\operatorname{Sg}({\mathbf{a}}^{1},\dotsc,{\mathbf{a}}^{n})\cap B^{n}=\emptyset$. Put $R=\operatorname{Sg}({\mathbf{a}}^{1},\dotsc,{\mathbf{a}}^{n})$ and assume that $R$ is an inclusion minimal relation among all choices of ${\mathbf{a}}^{1},\dotsc,{\mathbf{a}}^{n}\in A^{n}$.
@@
$\operatorname{Sg}\left((\{a^{n}_{n}\}\times B)\cup(B\times\{a^{n}_{n}\})\right).$
@@
We pp-define $R^{\prime}\leq$ by the formula
@@
For $i\in\{1,\dots,n\}$ by ${\mathbf{c}}^{i}$ we denote ${\mathbf{a}}^{i}$ with the last coordinate removed. By the definition of $R^{\prime}$ and $S$, we have $({\mathbf{c}}^{i},{\mathbf{c}}^{n}),({\mathbf{c}}^{n},{\mathbf{c}}^{i})\in R^{\prime}$ for every $i\in\{1,\dots,n-1\}$. Moreover, these $22n-$ 2 tuples satisfy the condition in the second part of Proposition 7.2. Therefore, if $R^{\prime}\cap B^{2n-2}=\emptyset$, then $B\not\trianglelefteq_{2n-2}$. But $B\trianglelefteq_{n+1}$ and $12n-2\geq n+$ 1, a contradiction.
Since $R^{\prime}\cap B^{2n-2}\neq\emptyset$, there exist ${\mathbf{d}},{\mathbf{d}}^{\prime}\in R\cap(B^{n-1}\times A)$ such that $(d_{n},d^{\prime}_{n})\in S$. Let $E$ be the projection of $R$ onto the last coordinate after fixing all the other coordinates to $B$, that is, $E=\operatorname{proj}_{n}(R\cap(B^{n-1}\times A))$. Since $R$ was chosen inclusion minimal, we get $\operatorname{proj}_{n}(R)\subseteq\operatorname{Sg}(B\cup\{e\})$ for every $e\in E$, otherwise we could replace ${\mathbf{a}}^{n}$ by a tuple ${\mathbf{b}}\in R\cap(B^{n-1}\times\{e\})$ and get a $B$-essential relation properly contained in $R$.
Let $E^{\prime}=E+S$. Since $B\cup\{d^{\prime}_{n}\}\subseteq E^{\prime}$ (as $a^{n}_{n},d_{n}\in E$, $(d_{n},d^{\prime}_{n})\in S$, and $(a^{n}_{n},b)\in S$ for all $b\in B$), we have $E^{\prime}\supseteq\operatorname{proj}_{n}(R)$, in particular $a_{n}^{n}\in E^{\prime}$. Therefore $(e,a_{n}^{n})\in S$ for some $e\in E$. It follows that $E^{\prime\prime}=a_{n}^{n}-S$ contains $B\cup\{e\}$ and we get $E^{\prime\prime}\supseteq\operatorname{proj}_{n}(R)$, in particular $a_{n}^{n}\in E^{\prime\prime}$. So $(a^{n}_{n},a^{n}_{n})\in S$, which contradicts our assumption on $B$.
The argument in Corollary 8.6 together with Lemma 8.10 and Proposition 8.12 now proves Proposition 4.3 (which is Corollary 7.10.2 in [Zhuk20:dichotomy]) and also the improved absorption theorem.
@@
Proof 8.14.
The algebra has no nontrivial projective subuniverses by Proposition 8.2, and therefore it has a transitive term operation by Proposition 8.1. By Proposition 8.4, $B$ is an absorbing subuniverse of . Lemma 8.10 and Proposition 8.12 now imply $B\trianglelefteq_{3}$.
@@
Proof 8.15.
Proposition 4.1 implies that there is a left central subuniverse of or . In both cases, Proposition 4.3 finishes the proof.
@@
Proof 8.16.
Let $R\leq_{sd}$ be the witnessing relation from the definition of center. We define a directed graph on $A$, by putting $c\rightarrow d$ if there exists a cyclic term $t$ and a choice of elements $b_{2},\dotsc,b_{n}\in B$ such that $t(c,b_{2},\dotsc,b_{n})=d$.
Obviously the graph has no sinks, but it is also easy to see that it is transitively closed. Indeed, let $c\rightarrow c^{\prime}$ be witnessed by $t$ and $b_{2},\dotsc,b_{n}$, and $c^{\prime}\rightarrow c^{\prime\prime}$ be witnessed by $t^{\prime}$ and $b_{2}^{\prime},\dotsc,b_{n^{\prime}}^{\prime}$. Then the star composition of $t^{\prime}$ and $t$, together with the tuple
@@
Next we claim that there is no $a\in A\setminus B$ with a self-loop. For a contradiction, say $a$ is such and let $t(a,b_{2},\dotsc,b_{n})=a$ with $a\in A\setminus B$ and $b_{2},\dotsc,b_{n}\in B$. The set $D=a+R$ is a subuniverse of , and moreover
$t(a+R,b_{2}+R,\dotsc,b_{n}+R)=t(D,C,\dotsc,C)\subseteq D$
and the same holds for all the cyclic shifts of $t$. Therefore, $D$ is a $2$-absorbing subuniverse of (by the operation $t(x,y,\dots,y)$), a contradiction.
We conclude that all the loops and directed cycles in this graph must be inside $B$. Therefore there is a number $m$ so that starting with any $a\in A$, and following any directed walk, after $m$ steps we necessarily arrive in $B$. It suffices to take any cyclic term of and star-compose it with itself $m$ times to obtain an operation which witnesses . Now it is enough to apply Lemma 8.10 and Proposition 8.12
@@
It follows that $lR\subseteq 2lR\subseteq 3lR\subseteq\ldots$, therefore this sequence of proper relations eventually stabilizes. We define $n=kl$ for a sufficiently large $k$ and obtain $2nR=nR$.
@@
Thus $T_{j}$ is proper, subdirect, central and pp-definable from $R$. However, its left center is a proper superset of the left center of $R$ since it contains the element $a_{j}$ which is not in the left center of $R$. This is a contradiction with the choice of $R$.